# MUG: Pseudo Labeling Augmented Audio-Visual Mamba Network for Audio-Visual Video Parsing

### v1

@@ 4.1 Experimental setup
MUG+ - 69.8 70.5 63.1 67.8 68.5 63.3 67.0 56.8 62.4 61.4
Table1: Comparison with the state-of-the-art methods on the LLP dataset in terms of F-scores. MUG+ further uses CLIP and CLAP as feature extractors and significantly boosts all metrics.
@@ A V AV Type@AV Event@AV
Temporal-Spatial attention (TSA). After extracting the features, we implement a simple temporal-spatial attention mechanism by Mamba. This mechanism operates on the attention of ten frames or ten audio segments of each video, which effectively captures the salient features in a video and strengthens the feature representation of different segments. From the experimental results (Table2), it can be seen that all metrics show a decline when TSA is not used.
Adaptive mamba fusion (AMF). We introduce additional branch based on Vision Mamba [50] for dynamically adjusting the scanning order. As shown in Table2, we can see that our approach achieves improvement. In the AVVP task, the events occurring in the two modalities are random, potentially being complementary or unrelated. Therefore,the AMF module shares cross-modal information while ensuring the independence of unimodal information. Besides, the addition of dynamic ordering branches mitigates the forgetting problem that mamba models pose in causal learning, which ignores information in early segments. More comparative experiments will be shown in supplementary.
Mamba feature enhancement module (MFE). We design MFE to capture inter-modal similarities and enrich feature learning. The ablation experiment presented in Table2 demonstrates the effectiveness of our proposed module. The metrics of the model are slightly improved after using MFE. The module amplifies the similar features of the two modalities at the same segment, thus allowing the two modalities to achieve a complementary effect in the prediction. We consider that when events occurring in both modalities are similar, one modality can assist in predicting events in the other modality (e.g. hearing an engine and seeing a car). In the AVVP task, the similar features of two modalities have a greater probability of representing similar events, while the complementary features have a greater probability of representing different events. When the events occurring in the two modalities are not similar, noise tends to be introduced during modality interaction. MFE and AMF confirm the necessity to balance contribution in prediction by cross-modal interaction. More comparative experiments will be shown in supplementary.

### v2

@@ 4.1 Experimental setup
Table1: Comparison with the state-of-the-art methods on the LLP dataset in terms of F-scores.
@@ A V AV Type@AV Event@AV
Mamba-based attention (MBA). After extracting the features, we implement a simple segment-based attention mechanism by Mamba. This mechanism operates on the attention of ten frames or ten audio segments of each video, which effectively captures the salient features in a video and strengthens the feature representation of different segments. From the experimental results (Table2), it can be seen that all metrics show a decline when MBA is not used.
Adaptive mamba fusion (AMF). We introduce additional branch based on Vision Mamba [50, 42] for dynamically adjusting the scanning order. As shown in Table2, we can see that our approach achieves improvement. In the AVVP task, the events occurring in the two modalities are random, potentially being complementary or unrelated. Therefore,the AMF module shares cross-modal information while ensuring the independence of unimodal information. Besides, the addition of dynamic ordering branches mitigates the forgetting problem that mamba models pose in causal learning, which ignores information in early segments.
Mamba feature enhancement module (MFE). We design MFE to capture inter-modal similarities and enrich feature learning. The ablation experiment presented in Table2 demonstrates the effectiveness of our proposed module. The metrics of the model are slightly improved after using MFE. The module amplifies the similar features of the two modalities at the same segment, thus allowing the two modalities to achieve a complementary effect in the prediction. We consider that when events occurring in both modalities are similar, one modality can assist in predicting events in the other modality (e.g. hearing an engine and seeing a car). In the AVVP task, the similar features of two modalities have a greater probability of representing similar events, while the complementary features have a greater probability of representing different events. When the events occurring in the two modalities are not similar, noise tends to be introduced during modality interaction. MFE and AMF confirm the necessity to balance contribution in prediction by cross-modal interaction.